OpenStudy (anonymous):

Help with this problem 3x+4/x=7 I can get the problem to this 4=7x-3^2......4=x(7-3x) but I don't know where to go from here

OpenStudy (mertsj):

\[\frac{3x+4}{x}=\frac{7}{1}\]

OpenStudy (mertsj):

Cross multiply

OpenStudy (mertsj):

\[7x=3x+4\]

OpenStudy (mertsj):

\[4x=4\]

OpenStudy (mertsj):

\[x=1\]

OpenStudy (anonymous):

The book says there is two answers 1 and 4/3 how did they get the 4/3

OpenStudy (mertsj):

Is the problem correct the way I have posted it?

OpenStudy (anonymous):

maybe not

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

 This is the equation from the book. I had to draw it as I don't know how to use the equation function.

OpenStudy (mertsj):